if a person walks at 4 meter/hr he covers a certain distance if he walks at 9 meter/hr he covers 7.5 miles more. How distance did he actually cover?
1
Expert's answer
2019-11-11T15:54:28-0500

Let the time of motion of a person is "t", then


"d_1=v_1t\/2"

"d_2=v_2t\/2"

Since


"d_2=d_1+12070\:\rm m"

we obtain


"9\:\rm m/hr\times {\it t}/2=4\:\rm m/hr\times {\it t}/2+12070\: m"


So the time of motion


"t=2\times\frac{12070\:\rm m}{9\:\rm m/hr-4\:\rm m/hr}=4828\:\rm hr"

The total distance


"d=d_1+d_2"

"=4\:\rm m/hr\times4828/2\: hr+9\:\rm m/hr\times4828/2\: hr"

"=31382\: \rm m=19.5\: mi"